OpenStudy (anonymous):

Solve by completing the square x^2-20x+30=0

OpenStudy (kaisertheslayer):

using this would help you and remember the a=1, b=-20, and c=30

OpenStudy (anonymous):

\[x^2-2*x*10+10^2-10^2+30=0\]

OpenStudy (anonymous):

what do you do after that? @surjithayer

OpenStudy (anonymous):

do i used the quadratic formula to solve by completing squares? @KaiserTheSlayer

OpenStudy (kaisertheslayer):

yes

OpenStudy (kaisertheslayer):

so it would be x=18.36660026534076 and x= 1.63339973465925

OpenStudy (anonymous):

oh okay thank you

OpenStudy (kaisertheslayer):

yup

OpenStudy (chillout):

Solving by square completion yields:\[x²-20x=-30\]\[x²-20x + 100 = 70\]\[(x-10)²=70\] There you have it. Do you the step-by-step?

OpenStudy (anonymous):

no i don't

OpenStudy (anonymous):

\[(x-10)^2-100+30=0\] \[\left( x-10 \right)^2=70=\left( \sqrt{70} \right)^2\] \[x-10=\pm \sqrt{70}\] x=?

OpenStudy (anonymous):

where is the 100 coming from?

OpenStudy (chillout):

You take half of the x coefficient, square it and add to both sides.

OpenStudy (anonymous):

thats right! i sorta remember now

OpenStudy (chillout):

Don't forget to divide everything by whatever is multiplying the squared term ;). In this case, it's "1", so it won't matter.

OpenStudy (anonymous):

so then to figure out what x is I solve... x-10=\[\sqrt{70}\] and x-10=\[\sqrt{-70}\]

OpenStudy (anonymous):

idk why it looks like that i was trying to write what you did on the last thing

OpenStudy (chillout):

Be careful! The second solution you gave is a COMPLEX number.

OpenStudy (anonymous):

i know i have to do the whole i stuff for that cause its a negative square root

OpenStudy (chillout):

No, it's not:\[x-10=\sqrt{70} \]or\[x-10=-\sqrt{70}\]

OpenStudy (chillout):

notice the minus sign is outside.

OpenStudy (anonymous):

yeah i see that now

OpenStudy (chillout):

And that's it! you have your solutions! Just move the independent term to the other side.

OpenStudy (anonymous):

yeah i got it now, thank you!

OpenStudy (chillout):

The whole thing about completing the square is to make the left side equation look like something like (x-a)² or (x+a)²